2.P028.WI. A truck covers 30.0 m in 8.60 s while smoothly slowing down to final speed of 2.90 m/s. (a) Find its original speed m/s (b) Find its acceleration. (Indicate the direction of the acceleration with the sign of your answer.) m/s Show My Work (optona truck in a uniform manner in an additional 5.00 s.
Answer #1

la Sa **(u+ v)x+ 30 = 4* (u + 2.9) x 8.6 u=4.08 m/s Answer 1b) v= ut at 2.9 – 4.08 + ax 8-6 > a= Answer - 0.137 m/s²